Fill in the missing fractions.
 ▯ - 3 / 21 = 5 / 21 ?????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given▯- 3 / 21 = 5 / 21
▯= 5 / 21 + 3 / 21
▯= (5 + 3) / 21
▯= 8 / 21

Thanks